# Topology of Discrete Quantum Feedback Control

### v1

@@ VIII.1 Formalism
The difference from the case with unitary feedback is the exitense of the additional index $\alpha$, which can be regarded as information that cannot be specified in the feedback process.
VIII.2 Dissipative Maxwell’s demon
@@
where $\mathcal{L}_{m}$ is the generator (Lindbladian) conditioned on the measurement outcome, $H_{m}$ is a feedback Hamiltonian given in Sec. IV, and $\hat{L}_{m,j}$ is the Lindblad operator acting on a local region around site $j$. Then, a CPTP map for the feedback operation is given by
@@
which takes the same form as in the case with unitary feedback control [see Eq. (43)]. Thus, regardless of whether the feedback control is unitary or not, the eigenvalue of the quantum channel and its topology are completely determined from the probability distribution $p(j|m)$ of the position after the feedback.
@@
$L_{m,j}^{(\mathrm{R})}=$ $\sqrt{\gamma_{\mathrm{R}}}\ket{j+1}\bra{j}.$ | . (180)
@@
VIII.3 Feedback control with engineered dissipation
@@
IX Conclusion and outlook
In summary, we have presented a class of dynamical topological phases realized with feedback control, which are chracterized by topology of quantum channels. The nontrivial point-gap topology of a quantum channel chracterized by a winding number ensures robust chiral tranport, which is analogous to the chiral edge current in the quantum Hall effect, and leads to the emergence of the non-Hermitian skin effect under the open boundary condition. We have also provided a symmetry classification of quantum feedback control, identified ten classes compatible with projective measurement, and constructed examples of feedback control in each symmetry class. This general framework of symmetry and topology of quantum feedback control provides a guiding princple for designing stable feedback control of quantum systems since topology is insensitive to small deformation. Also, feedback control offers a unique way of control of dynamical topological phases of matter distinct from other well-known classes such as periodically driven systems [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], non-Hermitian systems [32, 33, 34, 35, 36, 37, 38], and open quantum systems subject to engineered dissipation [64, 65].
@@ IX Conclusion and outlook
We conclude this paper by discussing future perspectives and open problems. First, while we have focused on point-gap topology of quantum channels, line-gap topology remains to be investigated. In analogy with dissipative topological insulators [34], nontrivial line-gap topology of quantum channels may lead to boundary-localized states and it can therefore be used to create and stabilize topological boundary states with feedback control.
Second, while we have mostly focused on feedback control with projective measurement, the present framework for topology of quantum channels is also applicable to general quantum measurement, and therefore it is natural to ask whether non-projective measurement can be harnessed to construct topological feedback control. In particular, since the symmetry classification in Sec. V.3 is based on the spectral property of projective measurement channels, BL symmetry classes other than those in Table 2 are not prohibited for general quantum measurement. Since a quantum channel in a symmetry class beyond the ten-fold way necessarily has a pair of eigenvalues ±$\pm 1$ (if one assumes the existense of a steady state), its eigenvalue spectrum is reminiscent of that of dissipative time crystals [33, 116], where the eigenvalue $-1$ implies oscillating components under repeated application of a quantum channel.
Third, a crucial distinction between quantum channels and general non-Hermitian matrices is complete positivity and trace preservation (CPTP), which should be satisfied by any time evolution of the density matrix of a quantum system [47]. Although the topological classification of non-Hermitian matrices without the CPTP condition has been obtained [34, 35], it is unclear whether the CPTP condition alters the topological classification. For example, in Sec. VI.2, we have shown that the CPTP condition places a nontrivial constraint on the Bloch matrix with transposition symmetry, which turns out to be characterized by an integer topological invariant rather than a $\mathbb{Z}_{2}$ invariant. It merits further study to investigate the generality of this phenomenon. From a broader perspective, this problem may be related to a classification of time evolution of open quantum systems described by CPTP maps.
@@ Acknowledgements.
This work was supported by KAKENHI Grant No. JP20K14383 and No. JP22H01152 from the Japan Society for the Promotion of Science. We gratefully acknowledge the support from the CREST program “Quantum Frontiers” (Grant No. JPMJCR23I1) by the Japan Science and Technology Agency.
@@ A.2 Momentum-space representation
for arbitrary $\hat{\rho}$ with $\hat{T}_{\lambda}\ (\lambda=1,\cdots,d)$ being translation operators:
@@
Here, we note that the operator $\tilde{c}_{\vec{i},a,c,\vec{\mu}}^{\dagger}$ has two additional indices $c$ and $\vec{\mu}$ compared with the original basis |$\ket{\vec{i},a}$ of the Hilbert space since $\tilde{\mathcal{E}}$ acts on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. Physically, this corresponds to the fact that the density matrix has bra and ket degrees of freedom.
@@ A.3 Mode expansion
Note that $\langle\hat{\rho}_{n}^{L}(\vec{k}),\hat{\rho}_{n}^{R}(\vec{k})\rangle=1$.
@@ Appendix B Feedback control with projective measurement
Since the Bloch matrix has nonzero matrix elements only in one columun, its eigenvalues $\xi_{n}(\vec{k})$ are given by its diagonal entries as
$\xi_{0}(\vec{k})=$ $X_{0,0}(\vec{k})$
@@
$\xi_{0}(\vec{k})=\sum_{\vec{r}}p(\vec{r})e^{-i\vec{k}\cdot\vec{r}},$ (230)
where $p(\vec{r})=p(\vec{R}_{j}-\vec{R}_{j^{\prime}}):=p(\vec{j}|\vec{j}^{\prime})$. This result indicates that the nonzero eigenvalue of the CPTP map can be obtained from the characteristic function of the probability distribution $\{p(\vec{r})\}$ of the displacement of a particle during feedback control. We note that $\xi_{0}(\vec{k}=0)=\sum_{\vec{r}}p(\vec{r})=1$, which is consistent with the trace-preserving condition (223).
The Bloch matrix for the $D_{\mathrm{loc}}=1$ case is diagonalizable if $X_{0,0}(\vec{k})\neq 0$. If $X_{0,0}(\vec{k})=0$, it cannot be diagonalized except for the trivial case with $X(\vec{k})=0$. Thus, the point at which $X_{0,0}(\vec{k})=\xi_{0}(\vec{k})=0$ corresponds to an exceptional (non-diagonalizable) point [68]. Specifically, if $X_{0,0}(\vec{k})\neq 0$, the left eigenvector $v_{n}^{L}(\vec{k})$ and the right eigenvector $v_{n}^{R}(\vec{k})$ corresponding to the eigenvalue $\xi_{n}(\vec{k})$ are given by
$v_{0}^{L}(\vec{k})=$ $(1,0,\cdots,0)^{T},$ (231)
$v_{0}^{R}(\vec{k})=$ $(1,X_{1,0}(\vec{k})/X_{0,0}(\vec{k}),\cdots,X_{N_{\mathrm{cell}}-1,0}(\vec{k})/X_{0,0}(\vec{k}))^{T},$ (232)
$v_{n}^{L}(\vec{k})=$ $(-X_{n,0}^{*}(\vec{k})/X_{0,0}^{*}(\vec{k}),0,\cdots,0)^{T}+e_{n}\ (n\neq 0),$ (233)
$v_{n}^{R}(\vec{k})=$ $e_{n},\ (n\neq 0),$ (234)
where $e_{n}$ is a unit vector with $(e_{n})_{\mu}=\delta_{\mu,n}$. These eigenvectors satisfy the biorthogonal condition $(v_{m}^{L}(\vec{k}),v_{n}^{R}(\vec{k}))=\delta_{m,n}$. Substituting Eqs. (231)-(234) into the expansions (212)-(214), we obtain the density matrix after feedback control as
$\mathcal{E}(\hat{\rho}_{\mathrm{i}})=$ $\frac{1}{\sqrt{N_{\mathrm{cell}}}}\sum_{\vec{k}}\sum_{\vec{j}}\sum_{\vec{\mu}}\xi_{0}(\vec{k})c(\vec{k})e^{i\vec{k}\cdot\vec{R}_{\vec{j}}}(v_{n}^{R}(\vec{k}))_{\vec{\mu}}$
@@
$+\frac{1}{\sqrt{N_{\mathrm{cell}}}}\sum_{\vec{k}}\sum_{\vec{j}}\sum_{\vec{\mu}\neq 0}c(\vec{k})e^{i\vec{k}\cdot\vec{R}_{\vec{j}}}X_{0,\vec{\mu}}(\vec{k})$
@@ Appendix D Correspondence between symmetry of a time-evolution operator and that of its generator
In Markovian open quantum systems, a generator (Lindbladian) $\mathcal{L}$ of a CPTP map $\mathcal{E}=e^{\mathcal{L}t}$ for time evolution features the Gorini-Kossakowski-Sudarshan-Lindblad form [97]. The matrix representation of a CPTP map through the Choi-Jamiołkowski isomorphism (see Sec. A.1) is denoted by ~$\tilde{\mathcal{E}}=e^{\tilde{\mathcal{L}}t}$. Here, we note that a Lindbladian cannot be a general non-Hermitian matrix since the real part of its eigenvalues must be non-positive. Because of this constraint, only the following set of Bernard-LeClair symmetries are possible for a Lindbladian [65] (see also Refs. [66, 67]):
@@ Appendix E Symmetry-preserving boundary condition for a $\mathbb{Z}_{2}$ demon
In this appendix, we describe how to implement a symmetry-preserving boundary condition for the $\mathbb{Z}_{2}$ demon in Sec. VII.4. The transition probability amplitudes for a feedback control in class ${}^{\dagger}$ must satisfy Eqs. (110a)-(110c). Here we perform a feedback control near the boundary of the system to preserve these symmetry conditions. We consider a 1D lattice with $L$ unit cells under the OBC, and the site $j=1$ ($j=L$) denotes the left (right) boundary. The feedback control under the PBC satisfies the conditions (110a)-(110c). However, if we perform the same feedback under the OBC, the reflection of a particle at the boundary violates these conditions. A feedback protocol that preserves the symmetry at the boundary is given as follows. Let $m_{1}=(j_{1},a_{1})$ be a measurement outcome of the first measurement. If $j_{1}=1$ or $j_{1}=L$, we do nothing (i.e., the feedback Hamiltonian is set to zero: $\hat{H}_{1,a_{1}}=\hat{H}_{L,a_{1}}=0$). The transition amplitudes from the boundary sites therefore vanish. If $j_{1}\neq 1,L$, we perform the feedback control with the same Hamiltonian $\hat{H}_{j_{1},a_{1}}$ as that for the PBC case described in Sec. VII.4. Then, we perform the second projective measurement and obtain an outcome $m_{2}=(j_{2},a_{2})$. If $j_{2}=2,\cdots,L-1$, we perform the same additional feedback control as the PBC case shown in Eq. (154). If $j_{2}=1$ or $j_{2}=L$, we perform a feedback operation with the unitary operator
@@
A solution of Eq. (249) can be obtained similarly. For the parameters employed in the numerical simulation for Fig. 17, the conditions (251) and (252) are met, and therefore we set the phase factors as Eqs. (253)-(255) in Fig. 17.
It is worthwhile to note that conditions (248) and (249) may not have a solution. In that case, the symmetry of a quantum channel should be violated under the OBC and the symmetry-protected non-Hermitian skin effect is not ensured. However, if the symmetry is broken only near the boundary, the transport of a particle towards the boundary due to the skin effect remains partially for some time scale before a particle approaches the boundary, since the dynamics in the bulk is unaltered [33].

### v2

@@ IX.1 Formalism
The difference from the case with unitary feedback is the existence of the additional index $\alpha$, which can be regarded as information that cannot be specified in the feedback process.
IX.2 Dissipative Maxwell’s demon
@@
where $\mathcal{L}_{m}$ is the generator (Lindbladian) conditioned on the measurement outcome, $H_{m}$ is a feedback Hamiltonian given in Sec. IV.1, and $\hat{L}_{m,j}$ is the Lindblad operator acting on a local region around site $j$. Then, a CPTP map for the feedback operation is given by
@@
which takes the same form as in the case with unitary feedback control [see Eq. (43)]. Thus, regardless of whether the feedback control is unitary or not, the eigenvalues of the quantum channel and its topology are completely determined from the probability distribution $p(j|m)$ of the position after the feedback.
@@
$L_{m,j}^{(\mathrm{R})}=$ $\sqrt{\gamma_{\mathrm{R}}}\ket{j+1}\bra{j},$ (190)
@@
IX.3 Feedback control with engineered dissipation
@@
X Conclusion and outlook
In this paper, we have presented a class of dynamical topological phases realized with feedback control, which are chracterized by topology of quantum channels. The nontrivial point-gap topology of a quantum channel chracterized by a winding number ensures robust chiral tranport, which is analogous to the chiral edge current in the quantum Hall effect, and leads to the emergence of the non-Hermitian skin effect under the open boundary condition. We have also provided a symmetry classification of quantum feedback control, identified ten classes compatible with projective measurement, and constructed examples of feedback control in each symmetry class. This general framework of symmetry and topology of quantum feedback control provides a guiding princple for designing stable feedback control of quantum systems since topology is insensitive to small deformation. Also, feedback control offers a unique way of control of dynamical topological phases of matter distinct from other well-known classes such as periodically driven systems [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], non-Hermitian systems [32, 33, 34, 35, 36, 37, 38], and open quantum systems subject to engineered dissipation [64, 65].
@@ X Conclusion and outlook
We conclude this paper by discussing future perspectives and open problems. First, while we have mostly focused on feedback control with projective measurement, the present framework for topology of quantum channels is also applicable to general quantum measurement, and therefore it is natural to ask whether non-projective measurement can be harnessed to construct topological feedback control. In particular, since the symmetry classification in Sec. V.3 is based on the spectral property of projective measurement channels, BL symmetry classes other than those in Table 2 are not prohibited for general quantum measurement. Since a quantum channel in a symmetry class beyond the ten-fold way necessarily has a pair of eigenvalues ±$\pm 1$ (if one assumes the existence of a steady state), its eigenvalue spectrum is reminiscent of that of dissipative time crystals [33, 121], where the eigenvalue $-1$ implies oscillating components under repeated application of a quantum channel.
Second, as discussed in Sec. VIII, a crucial distinction between quantum channels and general non-Hermitian matrices is complete positivity and trace preservation (CPTP), which should be satisfied by any time evolution of the density matrix of a quantum system [47]. Although the topological classification of non-Hermitian matrices without the CPTP condition has been obtained [34, 35], it is unclear whether the CPTP condition alters the topological classification. For example, in Sec. VI.2, we have shown that the CPTP condition places a nontrivial constraint on the Bloch matrix with transposition symmetry, which turns out to be characterized by an integer topological invariant rather than a $\mathbb{Z}_{2}$ invariant. It merits further study to investigate the generality of this phenomenon. From a broader perspective, this problem may be related to a classification of time evolution of open quantum systems described by CPTP maps.
@@ Acknowledgements.
M.N. acknowledges helpful discussions with Yuto Ashida, Kohei Kawabata, Masatoshi Sato, Kenji Shimomura, and Ken Shiozaki during the workshop YITP-T-24-03 on “Recent Developments and Challenges in Topological Phases” held at the Yukawa Institute for Theoretical Physics. This work was supported by KAKENHI Grant No. JP20K14383, No. JP22H01152, and No. JP24K16989 from the Japan Society for the Promotion of Science. We gratefully acknowledge the support from the CREST program “Quantum Frontiers” (Grant No. JPMJCR23I1) by the Japan Science and Technology Agency.
@@ A.2 Momentum-space representation
for arbitrary $\hat{\rho}$ with $\hat{T}_{\lambda}\ (\lambda=1,\cdots,d)$ being a translation operator:
@@
Here, we note that we consider a single-particle system although Eq. (210) is written like the second-quantized form. We also note that the operator $\tilde{c}_{\vec{i},a,c,\vec{\mu}}^{\dagger}$ has two additional indices $c$ and $\vec{\mu}$ compared with the original basis |$\ket{\vec{i},a}$ of the Hilbert space since $\tilde{\mathcal{E}}$ acts on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. Physically, this corresponds to the fact that the density matrix has bra and ket degrees of freedom.
@@ A.3 Mode expansion
Note that $\langle\hat{\rho}_{m}^{L}(\vec{k}),\hat{\rho}_{n}^{R}(\vec{k})\rangle=\delta_{mn}$.
@@ Appendix B Feedback control with projective measurement
Since the Bloch matrix has nonzero matrix elements only in one columun, its eigenvalues $\xi_{\vec{\mu}}(\vec{k})$ are given by its diagonal entries as
$\xi_{\vec{0}}(\vec{k})=$ $X_{\vec{0},\vec{0}}(\vec{k})$
@@
$\xi_{\vec{0}}(\vec{k})=\sum_{\vec{r}}p(\vec{r})e^{-i\vec{k}\cdot\vec{r}},$ (240)
where $p(\vec{r})=p(\vec{R}_{j}-\vec{R}_{j^{\prime}}):=p(\vec{j}|\vec{j}^{\prime})$. This result indicates that the nonzero eigenvalue of the CPTP map can be obtained from the characteristic function of the probability distribution $\{p(\vec{r})\}$ of the displacement of a particle during feedback control. We note that $\xi_{\vec{0}}(\vec{k}=0)=\sum_{\vec{r}}p(\vec{r})=1$, which is consistent with the trace-preserving condition (233).
The Bloch matrix for the $D_{\mathrm{loc}}=1$ case is diagonalizable if $X_{\vec{0},\vec{0}}(\vec{k})\neq 0$. If $X_{\vec{0},\vec{0}}(\vec{k})=0$, it cannot be diagonalized except for the trivial case with $X(\vec{k})=0$. Thus, the point at which $X_{\vec{0},\vec{0}}(\vec{k})=\xi_{\vec{0}}(\vec{k})=0$ corresponds to an exceptional (non-diagonalizable) point [68]. Specifically, if $X_{\vec{0},\vec{0}}(\vec{k})\neq 0$, the left eigenvector $v_{\vec{\nu}}^{L}(\vec{k})$ and the right eigenvector $v_{\vec{\nu}}^{R}(\vec{k})$ corresponding to the eigenvalue $\xi_{\vec{\nu}}(\vec{k})$ are given by
$(v_{\vec{0}}^{L}(\vec{k}))_{\vec{\mu}}=$ $\delta_{\vec{\mu},\vec{0}},$ (241)
$(v_{\vec{0}}^{R}(\vec{k}))_{\vec{\mu}}=$ $X_{\vec{\mu},\vec{0}}(\vec{k})/X_{\vec{0},\vec{0}}(\vec{k}),$ (242)
$(v_{\vec{\nu}}^{L}(\vec{k}))_{\vec{\mu}}=$ $-\delta_{\vec{\mu},\vec{0}}X_{\vec{\nu},\vec{0}}^{*}(\vec{k})/X_{\vec{0},\vec{0}}^{*}(\vec{k})+\delta_{\vec{\mu},\vec{\nu}}\ (\vec{\nu}\neq\vec{0}),$ (243)
$(v_{\vec{\nu}}^{R}(\vec{k}))_{\vec{\mu}}=$ $\delta_{\vec{\mu},\vec{\nu}}\ (\vec{\nu}\neq\vec{0}).$ (244)
These eigenvectors satisfy the biorthogonal condition $(v_{\vec{\mu}}^{L}(\vec{k}),v_{\vec{\nu}}^{R}(\vec{k}))=\delta_{\vec{\mu},\vec{\nu}}$. Substituting Eqs. (241)-(244) into the expansions (222)-(224), we obtain the density matrix after feedback control as
$\mathcal{E}(\hat{\rho}_{\mathrm{i}})=$ $\frac{1}{\sqrt{N_{\mathrm{cell}}}}\sum_{\vec{k}}\sum_{\vec{j}}\sum_{\vec{\mu}}\xi_{0}(\vec{k})c(\vec{k})e^{i\vec{k}\cdot\vec{R}_{\vec{j}}}(v_{\vec{0}}^{R}(\vec{k}))_{\vec{\mu}}$
@@
$+\frac{1}{\sqrt{N_{\mathrm{cell}}}}\sum_{\vec{k}}\sum_{\vec{j}}\sum_{\vec{\mu}\neq 0}X_{\vec{\mu},0}(\vec{k})c(\vec{k})e^{i\vec{k}\cdot\vec{R}_{\vec{j}}}$
@@ Appendix D Correspondence between symmetry of a time-evolution operator and that of its generator
In Markovian open quantum systems, the generator (Lindbladian) $\mathcal{L}$ of a CPTP map $\mathcal{E}=e^{\mathcal{L}t}$ for time evolution features the Gorini-Kossakowski-Sudarshan-Lindblad form [99]. The matrix representation of a CPTP map through the Choi-Jamiołkowski isomorphism (see Sec. A.1) is denoted by ~$\tilde{\mathcal{E}}=e^{\tilde{\mathcal{L}}t}$. Here, we note that a Lindbladian cannot be a general non-Hermitian matrix since the real part of its eigenvalues must be non-positive. Because of this constraint, only the following set of Bernard-LeClair symmetries are possible for a Lindbladian [65] (see also Refs. [66, 67]):
@@ Appendix E Symmetry-preserving boundary condition for a $\mathbb{Z}_{2}$ demon
In this appendix, we describe how to implement a symmetry-preserving boundary condition for the $\mathbb{Z}_{2}$ demon in Sec. VII.4. The transition probability amplitudes for a feedback control in class AII must satisfy Eqs. (111a)-(111c). Here we perform a feedback control near the boundary of the system to preserve these symmetry conditions. We consider a 1D lattice with $L$ unit cells under the OBC, and the site $j=1$ ($j=L$) denotes the left (right) boundary. The feedback control under the PBC satisfies the conditions (111a)-(111c). However, if we perform the same feedback under the OBC, the reflection of a particle at the boundary violates these conditions. A feedback protocol that preserves the symmetry at the boundary is given as follows. Let $m_{1}=(j_{1},a_{1})$ be a measurement outcome of the first measurement. If $j_{1}=1$ or $j_{1}=L$, we do nothing (i.e., the feedback Hamiltonian is set to zero: $\hat{H}_{1,a_{1}}=\hat{H}_{L,a_{1}}=0$). The transition amplitudes from the boundary sites therefore vanish. If $j_{1}\neq 1,L$, we perform the feedback control with the same Hamiltonian $\hat{H}_{j_{1},a_{1}}$ as that for the PBC case described in Sec. VII.4. Then, we perform the second projective measurement and obtain an outcome $m_{2}=(j_{2},a_{2})$. If $j_{2}=2,\cdots,L-1$, we perform the same additional feedback control as the PBC case shown in Eq. (155). If $j_{2}=1$ or $j_{2}=L$, we perform a feedback operation with the unitary operator
@@
A solution of Eq. (259) can be obtained similarly. For the parameters employed in the numerical simulation in Fig. 18, the conditions (261) and (262) are met, and therefore we set the phase factors as Eqs. (263)-(265) in Fig. 18.
It is worthwhile to note that conditions (258) and (259) may not have a solution. In that case, the symmetry of a quantum channel should be violated under the OBC and the symmetry-protected non-Hermitian skin effect is not ensured. However, if the symmetry is broken only near the boundary, the transport of a particle towards the boundary due to the skin effect is partially maintained for a certain time before a particle approaches the boundary, since the dynamics in the bulk is unaltered [33].
Figure 23: Schematic illustration of topological feedback control that shows the non-Hermitian skin effect but with no boundary-localized steady state. A two-level atom with internal states $g,e$ on a 1D lattice is considered. The excited state $e$ spontaneously decays into the ground state $g$ with rate $\gamma$. If a projective measurement finds the atom in the excited state, a potential wall with height $V$ is introduced on the left site to the particle.
Figure 24: Eigenspectrum of a quantum channel in Eq. (268). The parameters are set to be $J_{e}=1,J_{g}=0.2,V=10,\gamma=2$, and $\tau=1$. Blue curves and blue dots at the origin constitue the eigenspectrum under the PBC for infinite system size, and orange dots show the eigenspectrum under the OBC for the system size $L=20$.
Appendix F Topological feedback control that shows the non-Hermitian skin effect but with no boundary-localized steady state
The models of topological feedback control presented in Sec. IV, VII.1, VII.2, VII.3, VII.4, and IX show the non-Hermitian skin effects and boundary-localized steady states due to nontrivial point-gap topology of quantum channels. While the former is widely observed in non-Hermitian topological phases [33, 72, 73, 88, 38, 74], the latter does not exist in conventional non-Hermitian systems and thus is unique to topological quantum channels which preserve the positivity and trace of a density matrix (see Sec. VIII.1). Here, to distinguish these two topological phenomena, we present a model that shows the non-Hermitian skin effect but with no boundary-localized steady states. We consider a two-level atom on a 1D lattice with $L$ sites. The internal states of the atom consist of a ground state ($g$) and an excited state ($e$). We perform a projective measurement of the position and the internal state of the atom described by the measurement operator $\hat{M}_{j,a}=\ket{j,a}\bra{j,a}$ |, where |$\ket{j,a}$ denotes a quantum state of the atom at site $j=1,\cdots,L$ and internal state $a=g,e$. Then, depending on the measurement outcome $m=(j,a)$, we let the atom undergo the time evolution under the Lindblad equation
$\frac{d\hat{\rho}}{dt}=$ $\mathcal{L}_{m}(\hat{\rho})$
$=$ $-i[\hat{H}_{j,a},\hat{\rho}]+\sum_{i}\Bigl{(}\hat{L}_{i}\hat{\rho}\hat{L}_{i}^{\dagger}-\frac{1}{2}\{\hat{L}_{i}^{\dagger}\hat{L}_{i},\hat{\rho}\}\Bigr{)},$ (266)
where
$\hat{H}_{j,a}=$ $-\sum_{i,a^{\prime}}J_{a^{\prime}}(\ket{i,a^{\prime}}\bra{i+1,a^{\prime}}+\ket{i+1,a^{\prime}}\bra{i,a^{\prime}})$
$+\delta_{a,e}V\ket{j-1,e}\bra{j-1,e}$ | (267)
is the feedback Hamiltonian with hopping amplitudes $J_{e},J_{g}\in\mathbb{R}$ and potential height $V>0$, and $\hat{L}_{i}=\sqrt{\gamma}\ket{i,g}\bra{i,e}$ | is a Lindblad operator that describes spontaneous decay with rate $\gamma>0$ from the excited state to the ground state. That is, we impose a potential wall only if the internal state of the atom is detected to be the excited state. See Fig. 23 for a schematic illustration. The quantum channel for the feedback control is thus given by (see also Sec. IX)
$\mathcal{E}(\hat{\rho})=\sum_{m}e^{\mathcal{L}_{m}\tau}(\hat{M}_{m}\hat{\rho}\hat{M}_{m}^{\dagger}),$ (268)
where $\tau$ is the duration of the feedback control.
Figure 24 shows the eigenspectrum of the quantum channel (268). Here, a salient feature is that the band in the PBC spectrum that involves the steady state (the unit eigenvalue) does not form a loop and thus has a vanishing winding number. Physically, this is because the atom is in the internal ground state in the steady state and thus is not localized near the edges of the system. Meanwhile, the band located away from the steady state exhibits a loop structure. This leads to the non-Hermitian skin effect of an atom in the internal excited state, which has a finite lifetime due to spontaneous decay. Thus, the non-Hermitian skin effect in topological feedback control is not always accompanied by a boundary-localized steady state.